### v1

# Enhancing Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems

Abstract
Machine translation focuses mainly on high-resource languages (HRLs), while low-resource languages (LRLs) like Taiwanese Hokkien are relatively under-explored. This study aims to address this gap by developing a dual translation model between Taiwanese Hokkien and both Traditional Mandarin Chinese and English. We employ a pre-trained LLaMA2-7B model specialized in Traditional Mandarin Chinese to leverage the orthographic similarities between Taiwanese Hokkien Han and Traditional Mandarin Chinese. Our comprehensive experiments involve translation tasks across various writing systems of Taiwanese Hokkien and between Taiwanese Hokkien and other HRLs. We find that the use of a limited monolingual corpus also further improve the model’s Taiwanese Hokkien capabilities. We then utilize our translation model to standardize all Taiwanese Hokkien writing systems into Hokkien Han, resulting in further performance improvements. Additionally, we introduce an evaluation method incorporating back-translation and GPT-4 to ensure reliable translation quality assessment even for LRLs. The study contributes to narrowing the resource gap for Taiwanese Hokkien and empirically investigates the advantages and limitations of pre-training and fine-tuning based on LLaMA 2.
Keywords: low-resource language, large language model, neural machine translation, Taiwanese Hokkien
[…]
Enhancing Hokkien Dual Translation by Exploring and Standardizing of Four Writing Systems
| Bo-Han Lu, Yi-Hsuan Lin, En-Shiun Annie Lee, Richard Tzong-Han Tsai††thanks: Corresponding author: thtsai@g.ncu.edu.tw |
| Department of Computer Science and Information Engineering, National Central University, Taiwan |
| Department of Computer Science, Faculty of Science, University of Toronto |
| Center for GIS, Research Center for Humanities and Social Sciences, Academia Sinica, Taiwan |
| {110522028, 109502543}@cc.ncu.edu.tw, |
| annie.lee@cs.toronto.edu, thtsai@g.ncu.edu.tw |
Abstract content
1. Introduction
Machine translation (MT), as a crucial subfield of natural language processing (NLP), serves a vital role in overcoming language barriers by translating more texts into the desired language. However, current MT systems predominantly cater to high-resource languages (HRLs), posing significant challenges for low-resource languages (LRLs). Specifically, Taiwanese Hokkien, which is mainly spoken in Taiwan, southern China and a number of countries in Southeast Asia (Ding, 2016a), faces unique issues owing to historical factors (Ding, 2016b) and a persistent absence of standardized writing systems. These factors lead to an extra layer of complexity by introducing inconsistent corpus hinders the development of NLP research and data-hungry translation models for this language.
[…]
With the recent advancement of large language models (LLMs) like BLOOM (Scao et al., 2022), ChatGPT and LLaMA (Touvron et al., 2023a), these models have demonstrated their capabilities across various multilingual NLP tasks, including translation tasks (Jiao et al., 2023; García et al., 2023; Yang et al., 2023; Xu et al., 2023). Despite these advancements, state-of-the-art LLMs leave room for improvement in translation tasks, particularly for languages that are considerably removed from HRL (Jiao et al., 2023; Hendy et al., 2023). This study employs a pre-trained LLaMA 2 (Touvron et al., 2023b) model specialized in Mandarin Chinese (ZH), aiming to leverage the orthographic similarities between HAN and ZH to develop a translation model capable of translating between different writing systems of Taiwanese Hokkien as well as between Taiwanese Hokkien and other HRLs like ZH and English.
| Language | Writing | Abbreviation | Script | Translated Example |
| Hokkien | Tâi-lô | (TL) | Latin | Tsit kuí kang lóng sī pháinn-thinn. |
[…]
| Hokkien | Hàn-lô | (HL) | *Hybrid | 這幾工 lóng 是歹天。 |
| Hokkien | Han | (HAN) | Chinese character | 這幾工攏是歹天。 |
| Mandarin | Chinese | (ZH) | Chinese character | 這幾天都是壞天氣。 |
| English | (EN) | Latin | These few days the weather has not been good. |
We conduct a comprehensive set of experiments involving translation between the Latin script and Chinese character writing systems of Taiwanese Hokkien as well as translation to and from ZH and English. Our findings indicate that the use of a monolingual corpus covering all Taiwanese Hokkien writing systems positively impacts the model’s dual translation performance. Contrary to expectations, extending the model’s vocabulary for Taiwanese Hokkien does not yield improvements in these capabilities. We also observe that incorporating parallel datasets involving HRL improves the model’s performance, while adding such datasets between the two different Taiwanese Hokkien scripts has detrimental effects.
We further tried to enhance the HANZH and HANEN translation by standardizing all Taiwanese Hokkien monolingual corpora into HAN before continued pre-training. The standardization procedure and training flow are illustrated in Figure 1. Experimental results suggest that this pre-processing step can slightly improve the average translation performance.
For reliable automatic evaluation of translation results, in addition to BLEU score (Papineni et al., 2002) and chrF++ (Popović, 2017) metrics, we modify Kocmi and Federmann (2023)’s evaluation prompt and incorporates a back-translation method so that GPT-4 (OpenAI, 2023) can make reliable evaluations even if the target language is a LRL.
We plan to release a translation system that includes HAN, POJ, ZH and English. We anticipate that this will serve as a reliable translation tool for the public community, and foster the generation of diverse datasets for Taiwanese Hokkien.
To summarize, the major contributions of this work are:
1.
Develop and release the first dual translation model for Taiwanese Hokkien, thereby narrowing the resource gap for this low-resource language.
2.
Empirical evidence to support the enhancement of model performance through monolingual corpora on top of parallel data.
3.
Standardized all Taiwanese Hokkien monolingual corpora into HAN prior to continued pre-training, leading to performance enhancements in translations among ZH, English, and HAN.
4.
Introduction of back-translation of LRL into HRL for GPT prompt-based evaluation.
2. Background
Taiwanese Hokkien, also referred to Hokkien, Hoklo, Taigi, Southern Min, or Min Nan, is a unique subset of the Southern Min dialects. Sharing a common linguistic heritage with the Fujian dialects, Taiwanese Hokkien has undergone an independent evolution influenced by a number of external factors, including indigenous languages and the colonial legacies of the Dutch and Japanese (Liao et al., 2020). In the following discussion, this dialect will be referred to as Hokkien for the purpose of simplicity.
Although Hokkien ranks as the second most common language spoken in Taiwan, recent census222https://www.stat.gov.tw/public/Data/
1112144316VT5YTOVB.pdf suggest a looming crisis due to the decreasing proficiency among younger generations. The challenge is compounded by the fact that most Hokkien speakers only have oral proficiency in Hokkien and lack familiarity with written forms. This dual challenge of decreasing oral proficiency and limited literacy underscores the urgency for targeted research.
2.1. Writing System Diversity in Hokkien
The writing systems of Hokkien can be divided into three main groups. The first is Hokkien Han (HAN), which is based on Traditional Chinese with additional characters, followed by Latin script systems such as Tâi-lô (TL) and Pe̍h-ōe-jī (POJ). In addition, a hybrid system known as Hàn-lô (HL) combines elements of both systems. Since 2009, an official orthography for HAN has been established and is currently used in educational systems. However, due to its relatively recent standardization, corpus resources of HAN are scarce compared to other systems. On the other hand, POJ, which was introduced by missionaries in the 19th century, has a substantial amount of digitized historical texts and thus provides a rich corpus of Hokkien writings. Moreover, TL, an adaptation of POJ, maintains a systematic correspondence with it. HL, with its mixture of Latin script and Chinese characters, exhibits considerable variance across textual resources due to the lack of a uniform standard for determining the use of Latin script or Chinese characters. Therefore, in this study, HL corpus is considered for training purposes, but is excluded from translation evaluations.
2.2. Semantic Divergence of Shared Chinese Characters in HAN and ZH
Despite the commonality of Chinese characters between HAN and ZH, many homographs differ semantically. For example, the term “手指” translates to “finger” in ZH, while it means “ring” in HAN. Moreover, common HAN terms often correspond to rarely used Chinese characters in ZH, such as ‘覕’ (hide) and ‘啉’ (drink). As a result, training a reliable translation model capable of translating between these two languages under low-resource conditions still remains challenge.
3. Related Work
3.1. Large Language Models in Translation
LLMs have recently made remarkable progress in translation tasks due to their robust language understanding capabilities from pre-training on massive corpora. In the field of applying LLMs to translation tasks, Moslem et al. (2023); Lin et al. (2022); Zhu et al. (2023); Zhang et al. (2023a); Vilar et al. (2023); García et al. (2023) attempted to use in-context learning (ICL) (Brown et al., 2020) to enhance the translation capabilities of LLMs. Their study demonstrated how the pattern of in-context learning, the selection of few-shot sentences, and their quantity could impact the translation results. Zhang et al. (2023b); Yang et al. (2023); Li et al. (2023) tried to enhance the translation abilities of LLMs through instruction-tuning (Ouyang et al., 2022) with small amounts of parallel data. Li et al. (2023) demonstrated 3 BLEU score on average advancements of multilingual translation when compared to the ICL method. Some research (Hendy et al., 2023; Jiao et al., 2023) has indicated that LLMs may have limited translation abilities because their language skills are largely shaped by training in English-centered texts. The translation proficiency of LLMs is often significantly limited when translating languages that are not linguistically close to English (Li et al., 2023). As a result, Yang et al. (2023) and Li et al. (2023) have included different translation languages for monolingual training. Xu et al. (2023)’s latest findings also adopt monolingual training before fine-tuning translation task. Using medium-sized models with 7B and 13B parameters, they surpassed GPT-3.5 and NLLB-54B (Team et al., 2022) in various translation tasks.
3.2. Neural Machine Translation in Hokkien
Due to the scarcity of training data, neural machine translation (NMT) for low-resource languages such as Hokkien faces unique challenges. Liao et al. (2022) has compiled a dataset for Hokkien speech recognition, which not only contributes to speech-related research, but also benefits NMT through the transcribed parallel data. The techniques of transfer learning and cross-lingual models offer possible ways to improve Hokkien NMT systems. Lu et al. (2022) investigates translation task between ZH, Hokkien code-mixing language and ZH. They apply transfer learning to utilize the knowledge pre-trained on ZH by XLM (Conneau and Lample, 2019), and develop a method to synthesize a code-mixing translation parallel dataset to achieve better translation results between code-mixing language and ZH. To the best of our knowledge, we are the first to explore the application of large language models to dual translation for Hokkien while considering both Latin script and Chinese character writing systems of Hokkien.
4. Methodology
4.1. Corpus Preparation
4.1.1. Monolingual Datasets
As for our continued pre-training data, we have gathered a wide range of linguistic resources from diverse sources that reflect the depth and diversity of the Hokkien language. We have included a comprehensive explanation of the dataset’s domain, writing system, and other essential characteristics in Table 2 and Figure 2.
Our dataset primarily comprises web articles in Hokkien collected from diverse internet sources including scripts from Hokkien recitation contest333http://ip194097.ntcu.edu.tw/longthok/longthok.asp, lyrics of Hokkien songs shared in Facebook communities444https://www.facebook.com/groups/922800454445724, and web-scraped articles covering various domains. The dataset also incorporates religious articles in Hokkien, content from Wikipedia, and Hokkien elementary school textbooks.
Additionally, we include subtitles from Hokkien television programs. Given that these subtitles often lack paragraph breaks and punctuation, we employed GPT-3.5-turbo to refine the textual structure, rendering it more akin to standard articles.
| Data Source | Word Count (M) | Sentence Count (K) | ||||
| HAN | HL | POJ | HAN | HL | POJ | |
| Recitation Contest | 0.36 | - | - | 38.1 | - | - |
| Hokkien Song Lyrics | 2.61 | - | - | 349.1 | - | - |
| Web-scraped Articles | 0.37 | 1.18 | 0.48 | 36 | 110 | 40.6 |
| Religious Texts | - | 2.48 | 3.44 | - | 299.5 | 412.9 |
| Wikipedia Articles | - | - | 1.34 | - | - | 126.2 |
| Hokkien Textbooks | 0.02 | - | - | 3.3 | - | - |
| Subtitles | 1.87 | - | - | 218.3 | - | - |
| Total | 5.23 | 3.66 | 5.26 | 644.8 | 409.5 | 579.7 |
| Data Source | Word Count (K) | Pairs | Language |
| Dictionary | 135.4 | 15195 | ZH EN Hokkien (POJ, HAN) |
| Technical Terms | 7.9 | 2677 | ZH EN Hokkien (POJ, HAN) |
| Religious Texts | 1090.7 | 17872 | Hokkien (POJ, HL) |
4.1.2. Parallel Datasets
For fine-tuning data, we collected the datasets from example sentences in the Hokkien dictionary555https://sutian.moe.edu.tw/zh-hant/, religious texts666http://pojbh.lib.ntnu.edu.tw/script/index.php, and technical terms777https://stti.moe.edu.tw/?lang=sutgi. For those texts that have corresponding parallel sentences in ZH, we utilized GPT-3.5-turbo to translate them into English. Further details regarding the writing system and other fundamental features for the datasets are provided in Table 3.
4.2. Model Training
Pre-trained Large Language Model
To leverage the shared Chinese character system between ZH and HAN, we chose TAIDE-7B888https://taide.tw/ as our base model. Enriched with an additional 24k Chinese character tokens and pre-trained on a 1.7B-token Traditional Chinese corpus, TAIDE-7B serves as an extension of LLaMA 2 (Touvron et al., 2023a), enhanced comprehension of Taiwan-specific Traditional Chinese terms.
Vocabulary Extension
Although the TAIDE-7B model contains a large amount of Chinese characters, it still lacks coverage for Hokkien Latin scripts and some rarely used Chinese characters specific to Hokkien. To address this, we further extend the vocabulary by training a sentence-piece (Kudo and Richardson, 2018) tokenizer on monolingual Hokkien corpora and merge it back to the original one. Specifically, the vocabulary was extended with an additional 130 Chinese character tokens for HAN and 1876 Latin script tokens for POJ, resulting in a final vocabulary size of 58,505.
Continued Pre-training
We conducted continued pre-training on monolingual Hokkien corpora across all writing systems. We followed procedures from Chinese-LLaMA-2 Cui et al. (2023) using Low-Rank Adaptation (LoRA) Hu et al. (2021) with gradient checkpoint to reduce computational cost, and trained 18 epochs to avoid undesirable outputs. Given the rule-based transformation between the POJ and TL systems, we converted all TL in the corpus to POJ using an existing tool999https://github.com/i3thuan5/KeSi, streamlining the model training across these two Latin script writing systems.
Translation Fine-tuning with Instruction
In the fine-tuning stage, we modified the LLaMA 2 instruction tuning template, using:
The label [TRANS] denotes translation, X and Y are the source and target sentences, respectively. The label [target_lang] indicates the target language for the translation. During fine-tuning, each language pair in the parallel data was fixed to 17,872 instances, including HAN-ZH, HAN-EN, POJ-ZH, POJ-EN, and POJ-HL, and each model was trained for one epoch.
Pre-training Corpus Script-Standardization
Given the increasing prevalence of HAN in Taiwanese communities in recent years, and aiming to better leverage the orthographic similarities between ZH and HAN, we consequently explored whether standardizing all Hokkien monolingual data into HAN could further improve translation performance in HANZH and HANEN directions. To achieve this, we employed our translation model that exhibited the best performance in POJ-HAN translations to standardize all Latin scripts in monolingual Hokkien data into the Chinese character system. Continued pre-training was then carried out on this standardized data.
4.3. Experimental Settings
4.3.1. Two Translation Testing Datasets
We used iCorpus101010https://github.com/Taiwanese-Corpus/icorpus_ka1_han3-ji7 as our testing data. It comprises news articles from various domains and includes both HAN and POJ. However under human evaluation, the HAN section of iCorpus contains a considerable number of lexical inaccuracies. To cover this issue in the test set, we first selected terms in HAN that deviate significantly from ZH, using their frequency as a selection criterion to ensure the translation difficulty. Based on this criterion, we sampled the top 100 sentences where HAN appears frequently and manually corrected their lexical according to official orthography. This resulted in a subset we named iCorpus-100. Due to the size limitation of iCorpus-100, we integrated an additional data from TAT (Liao et al., 2020), specifically to enhance the evaluation of ZHHAN and ENHAN translations. TAT comprises 2,661 parallel sentences sourced from the Taiwanese Across Taiwan speech recognition competitions. Using the same process as the training set, the EN side of the parallel data were generated through translation from ZH using GPT-3.5-turbo. Neither of these test data was used to continued pre-training and fine-tuning.
4.3.2. Evaluation Metrics
In order to conduct a comprehensive evaluation of our translation models, we used four different metrics: BLEU score (Papineni et al., 2002), chrF++ (Popović, 2017), and two additional GPT-based metrics. Unlike BLEU, chrF++ includes evaluations at the character level. This lead to significant difference between these two metrics when the target language is Latin script, especially POJ. We attribute this to the accent variations in POJ that may alter one or two characters within a word. Such accent-induced variations were not normalized in out corpus. As a result, the model may produce accent inconsistencies with the ground truth, leading to significant penalties in BLEU score, but relatively small drop in chrF++.
Though widely used in translation assessment, BLEU and chrF++ mainly focus on lexicon-level granularity and do not provide a comprehensive evaluation of overall translation quality. In particular, these measures are limited when applied to large language models, which often require more nuanced methods of evaluation (Hendy et al., 2023). Kocmi and Federmann (2023) indicates that GPT-4 holds promise for accurate evaluation of the translation outputs. Therefore, we employ GPT-4 to assess the generated translations, ranging from 0 to 100, by providing both the model’s translation and a reference. However, GPT-4’s ability to comprehend Hokkien is restricted, therefore to overcome this constraint, we implement back-translation for the model’s output when translating from ZH or English into Hokkien. The back-translated output is then compared with the source sentence for evaluation. This method is not suitable for scenarios where both the source and target sentences are in Hokkien. For the evaluation of translation scores from GPT-4, we conducted a human qualitative analysis on 20 samples from each of five score intervals, totaling 100 examples. The qualitative findings based on these analyses are summarized in 4.
After the evaluation with GPT-4, we calculate the GPT-4 score by taking the average of each translation result. Since translations scoring 80 or above closely approximate the original sentence’s meaning, we consider them as correct translations, which are used to compute the translation accuracy, named as GPT-4 accuracy.
| Score Interval | Quality Descriptors |
80-100 |
The translation matches the source in meaning, using identical terms or appropriate synonyms. The overall grammar is correct, and the meaning flows smoothly. |
60-79 |
Errors occur with the translation of proper nouns (e.g., names) or misuse of nouns or verbs, causing slight shifts in meaning. At times, some information may be omitted in the translation or there may be over-translations. |
40-59 |
Parts of the translation are unrelated to the source text. Misuse of verbs and nouns in inappropriate contexts occurs more frequently. |
20-39 |
Most of the translation has deviations from the source text, though occasionally it still conveys the basic idea. Grammar and sentence structures are often incorrect, making the translated content challenging to understand. |
0-19 |
The translation is almost entirely unrelated to the source text. Sentences are not fluent, with frequent major grammatical and structural errors. |
BLEU | chrF++ |
BLEU | chrF++ |
BLEU | chrF++ |
| ZH-HAN | HAN-ZH | Average | ||||||||||||||||||||||||
| BASELINE | 26.78 | 27.98 | 32.70 | 2 | 20.89 | 24.44 | 50.20 | 15 | 23.84 | 26.21 | 41.45 | 8.5 | ||||||||||||||
| NONE | 30.80 | 32.67 | 63.65 | 39 | 29.44 | 32.66 | 67.70 | 46 | 30.12 | 32.67 | 65.68 | 42.5 | ||||||||||||||
| HAN | 29.15 | 31.20 | 68.40 | 48 | 31.29 | 33.46 | 72.95 | 56 | 30.22 | 32.33 | 70.68 | 52.0 | ||||||||||||||
| ALL | 31.37 | 32.89 | 72.35 | 55 | 32.60 | 34.91 | 75.35 | 63 | 31.99 | 33.90 | 73.85 | 59.0 | ||||||||||||||
| ALL (EXT_VOCAB) | 30.90 | 32.43 | 70.30 | 49 | 33.02 | 34.94 | 74.90 | 58 | 31.96 | 33.69 | 72.60 | 53.5 | ||||||||||||||
| EN-HAN | HAN-EN | Average | ||||||||||||||||||||||||
| BASELINE | 6.72 | 13.15 | 48.70 | 16 | 12.32 | 39.88 | 58.25 | 25 | 9.52 | 26.52 | 53.48 | 20.5 | ||||||||||||||
| NONE | 11.34 | 17.75 | 60.30 | 25 | 14.39 | 41.77 | 58.60 | 21 | 12.87 | 29.76 | 59.45 | 23.0 | ||||||||||||||
| HAN | 12.83 | 18.72 | 64.50 | 37 | 16.15 | 42.87 | 64.55 | 36 | 14.49 | 30.80 | 64.53 | 36.5 | ||||||||||||||
| ALL | 12.82 | 18.62 | 65.80 | 40 | 18.45 | 44.92 | 66.75 | 38 | 15.64 | 31.77 | 66.28 | 39.0 | ||||||||||||||
| ALL (EXT_VOCAB) | 11.93 | 17.94 | 61.30 | 32 | 17.54 | 44.02 | 64.60 | 35 | 14.74 | 30.98 | 62.95 | 33.5 | ||||||||||||||
| ZH-POJ | POJ-ZH | Average | ||||||||||||||||||||||||
| BASELINE | 0.44 | 29.00 | 15.45 | 0 | 12.37 | 16.20 | 12.75 | 1 | 6.41 | 22.60 | 14.10 | 0.5 | ||||||||||||||
| NONE | 0.38 | 25.43 | 21.20 | 2 | 16.98 | 20.68 | 23.20 | 2 | 8.68 | 23.06 | 22.20 | 2.0 | ||||||||||||||
| HAN | 0.39 | 25.03 | 19.95 | 3 | 19.48 | 21.73 | 23.95 | 3 | 9.94 | 23.38 | 21.95 | 3.0 | ||||||||||||||
| ALL | 0.64 | 26.64 | 35.20 | 9 | 30.08 | 31.36 | 43.10 | 12 | 15.36 | 29.00 | 39.15 | 10.5 | ||||||||||||||
| ALL (EXT_VOCAB) | 0.61 | 28.83 | 38.05 | 9 | 28.15 | 29.85 | 36.00 | 7 | 14.38 | 29.34 | 37.03 | 8.0 | ||||||||||||||
| EN-POJ | POJ-EN | Average | ||||||||||||||||||||||||
| BASELINE | 0.16 | 16.21 | 6.25 | 0 | 1.94 | 21.43 | 4.90 | 0 | 1.05 | 18.82 | 5.58 | 0.0 | ||||||||||||||
| NONE | 0.19 | 17.75 | 11.05 | 0 | 3.35 | 23.69 | 11.30 | 0 | 1.77 | 20.72 | 11.18 | 0.0 | ||||||||||||||
| HAN | 0.21 | 17.52 | 8.45 | 0 | 2.92 | 24.36 | 9.25 | 0 | 1.57 | 20.94 | 8.85 | 0.0 | ||||||||||||||
| ALL | 0.19 | 19.89 | 20.05 | 1 | 7.83 | 30.53 | 24.20 | 3 | 4.01 | 25.21 | 22.13 | 2.0 | ||||||||||||||
| ALL (EXT_VOCAB) | 0.23 | 20.39 | 25.65 | 0 | 7.06 | 29.39 | 22.95 | 4 | 3.65 | 24.89 | 24.30 | 2.0 | ||||||||||||||
| HAN-POJ | POJ-HAN | Average | ||||||||||||||||||||||||
| BASELINE | 0.91 | 40.03 | - | - | 37.48 | 38.67 | - | - | 19.20 | 39.35 | - | - | ||||||||||||||
| NONE | 0.55 | 33.95 | - | - | 49.31 | 49.47 | - | - | 24.93 | 41.71 | - | - | ||||||||||||||
| HAN | 0.74 | 34.44 | - | - | 47.87 | 47.91 | - | - | 24.31 | 41.18 | - | - | ||||||||||||||
| ALL | 1.01 | 33.19 | - | - | 56.92 | 56.48 | - | - | 28.97 | 44.84 | - | - | ||||||||||||||
| ALL (EXT_VOCAB) | 0.93 | 36.40 | - | - | 52.13 | 51.93 | - | - | 26.53 | 44.17 | - | - | ||||||||||||||
| Fine-tuning Data | BLEU | chrF++ |
BLEU | chrF++ |
BLEU | chrF++ |
| ZH-HAN | HAN-ZH | Average | ||||||||||||||||||||||
| BASELINE | 18.70 | 21.98 | 65.70 | 48 | 14.88 | 19.15 | 63.85 | 48 | 16.79 | 20.57 | 64.78 | 48.0 | ||||||||||||
| ZH, HAN | 27.48 | 29.74 | 75.50 | 61 | 30.56 | 32.74 | 77.90 | 66 | 29.02 | 31.24 | 76.70 | 63.5 | ||||||||||||
| ZH, HAN, EN | 26.36 | 28.88 | 74.85 | 64 | 32.41 | 34.46 | 78.90 | 69 | 29.39 | 31.67 | 76.88 | 66.5 | ||||||||||||
| ZH, HAN, EN, POJ | 25.16 | 27.86 | 72.60 | 54 | 31.77 | 33.63 | 78.65 | 61 | 28.47 | 30.75 | 75.63 | 57.5 | ||||||||||||
| ZH, HAN, EN, POJ, HL | 31.37 | 32.89 | 72.35 | 55 | 32.60 | 34.91 | 75.35 | 63 | 31.99 | 33.90 | 73.85 | 59.0 | ||||||||||||
| EN-HAN | HAN-EN | Average | ||||||||||||||||||||||
| BASELINE | 9.49 | 15.04 | 41.35 | 14 | 2.62 | 22.51 | 44.05 | 15 | 6.06 | 18.78 | 42.70 | 14.5 | ||||||||||||
| ZH, HAN | 12.23 | 17.89 | 29.00 | 21 | 0.00 | 1.00 | - | - | 6.12 | 9.45 | 29.00 | 21.0 | ||||||||||||
| ZH, HAN, EN | 12.83 | 18.46 | 69.05 | 46 | 21.25 | 47.48 | 73.35 | 48 | 17.04 | 32.97 | 71.20 | 47.0 | ||||||||||||
| ZH, HAN, EN, POJ | 11.60 | 17.61 | 64.25 | 31 | 20.87 | 46.36 | 72.40 | 52 | 16.24 | 31.99 | 68.33 | 41.5 | ||||||||||||
| ZH, HAN, EN, POJ, HL | 12.82 | 18.62 | 65.80 | 40 | 18.45 | 44.92 | 66.75 | 38 | 15.64 | 31.77 | 66.28 | 39.0 | ||||||||||||
| ZH-POJ | POJ-ZH | Average | ||||||||||||||||||||||
| BASELINE | 0.08 | 3.36 | 18.30 | 7 | 1.33 | 4.13 | 2.80 | 0 | 0.71 | 3.75 | 10.55 | 3.5 | ||||||||||||
| ZH, HAN | 0.00 | 0.47 | - | - | 2.72 | 7.59 | 7.70 | 0 | 1.36 | 4.03 | 7.70 | 0.0 | ||||||||||||
| ZH, HAN, EN | 0.01 | 0.83 | - | - | 4.81 | 8.44 | 7.95 | 1 | 2.41 | 4.64 | 7.95 | 1.0 | ||||||||||||
| ZH, HAN, EN, POJ | 0.29 | 23.23 | 29.15 | 7 | 22.93 | 25.27 | 36.90 | 10 | 11.61 | 24.25 | 33.03 | 8.5 | ||||||||||||
| ZH, HAN, EN, POJ, HL | 0.64 | 26.64 | 35.20 | 9 | 30.08 | 31.36 | 43.10 | 12 | 15.36 | 29.00 | 39.15 | 10.5 | ||||||||||||
| EN-POJ | POJ-EN | Average | ||||||||||||||||||||||
| BASELINE | 0.02 | 4.70 | 2.50 | 0 | 0.20 | 6.85 | 1.40 | 0 | 0.11 | 5.78 | 1.95 | 0.0 | ||||||||||||
| ZH, HAN | 0.00 | 0.36 | - | - | 0.00 | 0.37 | 6.90 | 0 | 0.00 | 0.37 | 6.90 | 0.0 | ||||||||||||
| ZH, HAN, EN | 0.00 | 0.35 | - | - | 1.57 | 19.10 | 5.20 | 0 | 0.79 | 9.73 | 5.20 | 0.0 | ||||||||||||
| ZH, HAN, EN, POJ | 0.30 | 19.12 | 17.90 | 1 | 6.86 | 27.86 | 24.95 | 1 | 3.58 | 23.49 | 21.43 | 1.0 | ||||||||||||
| ZH, HAN, EN, POJ, HL | 0.19 | 19.89 | 20.05 | 1 | 7.83 | 30.53 | 24.20 | 3 | 4.01 | 25.21 | 22.13 | 2.0 | ||||||||||||
| HAN-POJ | POJ-HAN | Average | ||||||||||||||||||||||
| BASELINE | 0.04 | 4.37 | - | - | 1.01 | 4.13 | - | - | 0.53 | 4.25 | - | - | ||||||||||||
| ZH, HAN | 0.00 | 0.48 | - | - | 2.96 | 7.92 | - | - | 1.48 | 4.20 | - | - | ||||||||||||
| ZH, HAN, EN | 0.02 | 1.10 | - | - | 5.87 | 9.30 | - | - | 2.95 | 5.20 | - | - | ||||||||||||
| ZH, HAN, EN, POJ | 0.52 | 27.14 | - | - | 35.98 | 39.32 | - | - | 18.25 | 33.23 | - | - | ||||||||||||
| ZH, HAN, EN, POJ, HL | 1.01 | 33.19 | - | - | 56.92 | 56.48 | - | - | 28.97 | 44.84 | - | - | ||||||||||||
5. Experiment Results and Analysis
5.1. Experimental Ablation Studies
In order to isolate the impact of various data inputs, we conducted ablation studies on continued pre-training of three different Hokkien writing systems, the extension of the input vocabulary with Hokkien Chinese characters and Hokkien Latin systems, and fine-tuning these three different Hokkien writing systems111111Our focus on Hokkien LRL led us to exclude HRL data from training and evaluation..
5.1.1. Continued Pre-training Corpora and Vocabulary Extension Ablation Studies
As shown in Table 5, we took LLaMA 2-7B without any continued pre-training as a baseline. We compared its performance with the TAIDE-7B model continued pre-trained on different monolingual data and evaluated the impact of extending the Hokkien dictionary. Here, “NONE” indicates no continued pre-training on any Hokkien data, “HAN” indicates pre-training solely on Chinese character-based Hokkien data, and “ALL” indicates the inclusion of Latin script (POJ) and hybrid (HL) Hokkien data in addition to “HAN”.
The baseline model, which has no further continued pre-training, performs the worst. In contrast, the TAIDE-7B model improves significantly in all translation directions without relying on Hokkien monolingual data but ZH data. This suggests that using a similar HRL model as a foundational model is beneficial when supplementary monolingual data is not available. Pre-training on HAN data improves the GPT-4 score by 4 to 6 points in HAN-related translations. Incorporating all Hokkien data yields the best performance, particularly in POJ-related translations, with a 10 to 20 point increase in GPT-4 score. These findings align with previous research (Yang et al., 2023; Li et al., 2023; Xu et al., 2023), demonstrating the substantial performance improvement in translation tasks when monolingual data is utilized for languages that the foundational model is not familiar with.
Regarding vocabulary extension, since the added vocabulary primarily consists of Latin scripts from POJ, the model with vocabulary extension exhibit superior performance only when the target language is POJ. For other translation directions, these models exhibit a slight decrement, averaging 3 points lower on the GPT-4 score compared to models without vocabulary extension. We attribute this to the limited size of the pre-training corpus, which hinders effective tuning of newly added tokens. Consequently, we opted not to extend the vocabulary and suggest future work in collecting a larger POJ corpus for further investigation.
| Continued Pre-train Data | BLEU | chrF++ |
BLEU | chrF++ |
BLEU | chrF++ |
| ZH-HAN | HAN-ZH | Average | ||||||||||||||||||||||
| CP_HAN (w/o Standardized Data) | 30.63 | 32.16 | 70.58 | 54.38 | 35.20 | 37.55 | 77.96 | 68.92 | 32.92 | 34.86 | 74.27 | 61.65 | ||||||||||||
| CP_ALL | 32.16 | 34.90 | 74.21 | 62.53 | 35.17 | 37.03 | 77.87 | 67.98 | 33.67 | 35.97 | 76.04 | 65.26 | ||||||||||||
| CP_HAN (w/ Standardized Data) | 32.98 | 33.86 | 73.09 | 59.41 | 36.36 | 38.09 | 78.96 | 69.90 | 34.67 | 35.98 | 76.03 | 64.66 | ||||||||||||
| EN-HAN | HAN-EN | Average | ||||||||||||||||||||||
| CP_HAN (w/o Standardized Data) | 19.64 | 23.04 | 63.57 | 41.87 | 24.90 | 46.95 | 67.80 | 50.19 | 22.27 | 35.00 | 65.69 | 46.03 | ||||||||||||
| CP_ALL | 20.05 | 23.40 | 64.52 | 42.81 | 24.42 | 46.57 | 68.14 | 50.00 | 22.24 | 34.99 | 66.33 | 46.41 | ||||||||||||
| CP_HAN (w/ Standardized Data) | 20.99 | 24.31 | 67.32 | 47.05 | 24.36 | 46.90 | 69.12 | 52.31 | 22.68 | 35.61 | 68.22 | 49.68 | ||||||||||||
5.1.2. Fine-tuning Datasets Ablation Study
Given that the model using all Hokkien monolingual data without vocabulary extension performed the best on average across all translation directions, we selected it as the base model for further experiments in the fine-tuning stage. We investigated the impact of incorporating different parallel data during fine-tuning. The parallel data containing HL was only available in the HLPOJ direction. As a baseline, we followed the methodology of Li et al. (2023), utilizing in-context learning (ICL) for 8-shot translation without any fine-tuning.
Table 6 indicates that models fine-tuned on parallel datasets significantly outperform few-shot ICL models across all translation directions, particularly in directions involving Latin script Hokkien (POJ), with a GPT-4 score increase from 17 to 54 points. Demonstrate that the benefits of fine-tuning are particularly substantial for writing systems that are not closely related to HRL.
For ZHHAN and ENHAN directions, including ENHAN data enhance translation performance. However, further incorporation of POJ and HL parallel data does not yield additional improvements, suggesting that focusing solely on HRL in the parallel data is more effective for aligning cross-lingual embeddings.
In Hokkien script translation, the model performs the best when all parallel datas are included. The inclusion of POJHL data dramatically improve translation capabilities in the POJHAN. The chrF++ improved by 6.05 and 17.16 points for HAN-POJ and POJ-HAN directions, respectively. We hypothesize that HL contains a few Han characters, allowing the model to learn some lexical correlations between different scripts in Hokkien.
5.2. Pre-training Corpus Script-Standardization
We evaluated the efficacy of continued pre-training using three different datasets: 1) monolingual Hokkien data in Chinese characters only, 2) monolingual Hokkien data in all writing systems, and 3) all monolingual Hokkien data standardized into Han characters. These three models were subsequently fine-tuned using only ZHHAN and ENHAN parallel data.
Table 7 present the evaluation results of these three models on TAT. The model pre-trained on standardized monolingual data on par or slightly outperforms the other models on average across both pairs of languages. This suggests that standardizing the writing system into Chinese character can yield benefits on HAN translation.
6. Conclusion
We conducted a thorough examination of the efficiency of large language models in a translation system including ZH, English and two different Hokkien writing systems (HAN and POJ). We employed evaluation metrics including BLEU, chrF++, and a GPT-4-based scoring system. Our results showed that employing Hokkien similar high-resource language model as a foundational model led to significant performance enhancements. Furthermore, models pre-trained on all available Hokkien data exhibited the highest performance. Extending vocabulary appeared to favor the incorporation of more monolingual data to elicit their efficiency. In the fine-tuning stage, the model demonstrates enhanced HAN translation only when it is supplied with parallel data corresponding to the HRLs that were extensively involved in its prior pre-training. Additionally, we investigated the benefits of standardizing the Hokkien scripts to Chinese characters. Our future work will explore the potential advantage of data augmentation through translations from ZH monolingual corpora into Hokkien Han and assess its impact on translation quality. Moreover, extending this research to include other prevalent spoken languages in Taiwan, like Hakka, can offer a more extensive viewpoint on handling the linguistic variations in Taiwan.
7. Limitation
The methodology used in this study is conditional on the fact that one writing system in a low-resource language is similar to a high-resource language, which can be seen as a limitation. In this work, Hokkien Han and Mandarin Chinese share similar writing systems and possess a substantial amount of common vocabulary. This allows for a transfer of knowledge from extensive Mandarin Chinese texts, thereby leveraging the benefits of large language models pre-trained on abundant Mandarin Chinese corpora to achieve an exceptionally efficient translation system.
8. Ethical Considerations
One of the major ethical challenges in developing large language models for Hokkien is the limited resources and biased nature of available data. Most of the existing datasets come from news articles that exhibit specific ideological stances, political inclinations, or ethnic biases. The utilization of such skewed data may inadvertently train the model to propagate these biases, thus affecting its fairness. To address ethical concerns, we expanded our dataset to include lyrics, essays, and other neutral literary texts. Our goal was to reduce potential biases and create a more balanced and representative model.
9. Acknowledgements
We would like to express our sincere gratitude to Dr. Yu-Chun Wang for the generous assistance and insightful discussions on Taiwanese Hokkien. Our thanks also go to Ray Hung for his invaluable help in polishing the paper writing. Additionally, we are deeply grateful to all the reviewers for their dedication to improving this research. We are thankful to the Trustworthy AI Dialogue Engine (TAIDE) project for providing the Traditional Chinese foundation LLM, which served as a crucial base for building our model. Special appreciation goes to the National Center for High-performance Computing (NCHC) for providing computational and storage resources. This work was also supported in part by the Co-creation Platform of the Speech-AI Research Center, Industry-Academia Innovation School, NYCU, under the framework of the National Key Fields Industry-University Cooperation and Skilled Personnel Training Act, from the Ministry of Education (MOE), the National Development Fund (NDF), and industry partners in Taiwan.
10. Bibliographical References
[…]
- Conneau and Lample (2019) Alexis Conneau and Guillaume Lample. 2019. Cross-lingual language model pretraining. Advances in neural information processing systems, 32.
- Cui et al. (2023) Yiming Cui, Ziqing Yang, and Xin Yao. 2023. Efficient and effective text encoding for chinese llama and alpaca. arXiv preprint arXiv:2304.08177.
- Ding (2016a) Picus Sizhi Ding. 2016a. Introduction, pages 1–18. Springer Singapore, Singapore.
- Ding (2016b) Picus Sizhi Ding. 2016b. Taiwan: The Haven for Southern Min?, pages 55–75. Springer Singapore, Singapore.
- García et al. (2023) Xavier García, Yamini Bansal, Colin Cherry, George F. Foster, Maxim Krikun, Fan Feng, Melvin Johnson, and Orhan Firat. 2023. The unreasonable effectiveness of few-shot learning for machine translation. ArXiv, abs/2302.01398.
- Hendy et al. (2023) Amr Hendy, Mohamed Gomaa Abdelrehim, Amr Sharaf, Vikas Raunak, Mohamed Gabr, Hitokazu Matsushita, Young Jin Kim, Mohamed Afify, and Hany Hassan Awadalla. 2023. How good are gpt models at machine translation? a comprehensive evaluation. ArXiv, abs/2302.09210.
- Hu et al. (2021) Edward J. Hu, Yelong Shen, Phillip Wallis, Zeyuan Allen-Zhu, Yuanzhi Li, Shean Wang, Lu Wang, and Weizhu Chen. 2021. Lora: Low-rank adaptation of large language models.
- Jiao et al. (2023) Wenxiang Jiao, Wenxuan Wang, JT Huang, Xing Wang, and ZP Tu. 2023. Is chatgpt a good translator? yes with gpt-4 as the engine. arXiv preprint arXiv:2301.08745.
- Kocmi and Federmann (2023) Tom Kocmi and Christian Federmann. 2023. Large language models are state-of-the-art evaluators of translation quality. In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, pages 193–203, Tampere, Finland. European Association for Machine Translation.
- Kudo and Richardson (2018) Taku Kudo and John Richardson. 2018. SentencePiece: A simple and language independent subword tokenizer and detokenizer for neural text processing. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing: System Demonstrations, pages 66–71, Brussels, Belgium. Association for Computational Linguistics.
- Li et al. (2023) Jiahuan Li, Hao Zhou, Shujian Huang, Shan Chen, and Jiajun Chen. 2023. Eliciting the translation ability of large language models via multilingual finetuning with translation instructions. ArXiv, abs/2305.15083.
- Liao et al. (2020) Yuan-Fu Liao, Chia-Yu Chang, Hak-Khiam Tiun, Huang-Lan Su, Hui-Lu Khoo, Jane S. Tsay, Le-Kun Tan, Peter Kang, Tsun-guan Thiann, Un-Gian Iunn, Jyh-Her Yang, and Chih-Neng Liang. 2020. Formosa Speech Recognition Challenge 2020 and Taiwanese Across Taiwan Corpus. In 2020 23rd Conference of the Oriental COCOSDA International Committee for the Co-ordination and Standardisation of Speech Databases and Assessment Techniques (O-COCOSDA), pages 65–70.
- Liao et al. (2022) Yuan-Fu Liao, Jane S. Tsay, Peter Kang, Hui-Lu Khoo, Le-Kun Tan, Li-Chen Chang, Un-Gian Iunn, Huang-Lan Su, Tsun-Guan Thiann, Hak-Khiam Tiun, and Su-Lian Liao. 2022. Taiwanese across taiwan corpus and its applications. In 2022 25th Conference of the Oriental COCOSDA International Committee for the Co-ordination and Standardisation of Speech Databases and Assessment Techniques (O-COCOSDA), pages 1–5.
- Lin et al. (2022) Xi Victoria Lin, Todor Mihaylov, Mikel Artetxe, Tianlu Wang, Shuohui Chen, Daniel Simig, Myle Ott, Naman Goyal, Shruti Bhosale, Jingfei Du, Ramakanth Pasunuru, Sam Shleifer, Punit Singh Koura, Vishrav Chaudhary, Brian O’Horo, Jeff Wang, Luke Zettlemoyer, Zornitsa Kozareva, Mona Diab, Veselin Stoyanov, and Xian Li. 2022. Few-shot Learning with Multilingual Generative Language Models. In Proceedings of the 2022 Conference on Empirical Methods in Natural Language Processing, pages 9019–9052. Association for Computational Linguistics.
- Lu et al. (2022) Sin-En Lu, Bo-Han Lu, Chao-Yi Lu, and Richard Tzong-Han Tsai. 2022. Exploring methods for building dialects-Mandarin code-mixing corpora: A case study in Taiwanese hokkien. In Findings of the Association for Computational Linguistics: EMNLP 2022, pages 6287–6305, Abu Dhabi, United Arab Emirates. Association for Computational Linguistics.
- Moslem et al. (2023) Yasmin Moslem, Rejwanul Haque, John D. Kelleher, and Andy Way. 2023. Adaptive Machine Translation with Large Language Models. In Proceedings of the 24th Annual Conference of the European Association for Machine Translation, pages 227–237. European Association for Machine Translation.
- OpenAI (2023) OpenAI. 2023. Gpt-4 technical report. ArXiv, abs/2303.08774.
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback.
- Papineni et al. (2002) Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002. Bleu: a method for automatic evaluation of machine translation. In Proceedings of the 40th Annual Meeting of the Association for Computational Linguistics, pages 311–318, Philadelphia, Pennsylvania, USA. Association for Computational Linguistics.
- Popović (2017) Maja Popović. 2017. chrF++: words helping character n-grams. In Proceedings of the Second Conference on Machine Translation, pages 612–618, Copenhagen, Denmark. Association for Computational Linguistics.
- Scao et al. (2022) Teven Le Scao, Angela Fan, Christopher Akiki, Ellie Pavlick, Suzana Ilic, Daniel Hesslow, Roman Castagné, Alexandra Sasha Luccioni, François Yvon, Matthias Gallé, Jonathan Tow, Alexander M. Rush, Stella Biderman, Albert Webson, Pawan Sasanka Ammanamanchi, Thomas Wang, Benoît Sagot, Niklas Muennighoff, Albert Villanova del Moral, Olatunji Ruwase, Rachel Bawden, Stas Bekman, Angelina McMillan-Major, Iz Beltagy, Huu Nguyen, Lucile Saulnier, Samson Tan, Pedro Ortiz Suarez, Victor Sanh, Hugo Laurençon, Yacine Jernite, Julien Launay, Margaret Mitchell, Colin Raffel, Aaron Gokaslan, Adi Simhi, Aitor Soroa, Alham Fikri Aji, Amit Alfassy, Anna Rogers, Ariel Kreisberg Nitzav, Canwen Xu, Chenghao Mou, Chris Emezue, Christopher Klamm, Colin Leong, Daniel van Strien, David Ifeoluwa Adelani, and et al. 2022. BLOOM: A 176b-parameter open-access multilingual language model. CoRR, abs/2211.05100.
- Team et al. (2022) NLLB Team, Marta R. Costa-jussà, James Cross, Onur Çelebi, Maha Elbayad, Kenneth Heafield, Kevin Heffernan, Elahe Kalbassi, Janice Lam, Daniel Licht, Jean Maillard, Anna Sun, Skyler Wang, Guillaume Wenzek, Al Youngblood, Bapi Akula, Loic Barrault, Gabriel Mejia Gonzalez, Prangthip Hansanti, John Hoffman, Semarley Jarrett, Kaushik Ram Sadagopan, Dirk Rowe, Shannon Spruit, Chau Tran, Pierre Andrews, Necip Fazil Ayan, Shruti Bhosale, Sergey Edunov, Angela Fan, Cynthia Gao, Vedanuj Goswami, Francisco Guzmán, Philipp Koehn, Alexandre Mourachko, Christophe Ropers, Safiyyah Saleem, Holger Schwenk, and Jeff Wang. 2022. No language left behind: Scaling human-centered machine translation.
- Touvron et al. (2023a) Hugo Touvron, Thibaut Lavril, Gautier Izacard, Xavier Martinet, Marie-Anne Lachaux, Timothée Lacroix, Baptiste Rozière, Naman Goyal, Eric Hambro, Faisal Azhar, Aurelien Rodriguez, Armand Joulin, Edouard Grave, and Guillaume Lample. 2023a. Llama: Open and efficient foundation language models.
- Touvron et al. (2023b) Hugo Touvron, Louis Martin, Kevin Stone, Peter Albert, Amjad Almahairi, Yasmine Babaei, Nikolay Bashlykov, Soumya Batra, Prajjwal Bhargava, Shruti Bhosale, et al. 2023b. Llama 2: Open foundation and fine-tuned chat models. arXiv preprint arXiv:2307.09288.
- Vilar et al. (2023) David Vilar, Markus Freitag, Colin Cherry, Jiaming Luo, Viresh Ratnakar, and George Foster. 2023. Prompting PaLM for Translation: Assessing Strategies and Performance. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), pages 15406–15427. Association for Computational Linguistics.
- Xu et al. (2023) Haoran Xu, Young Jin Kim, Amr Sharaf, and Hany Hassan Awadalla. 2023. A paradigm shift in machine translation: Boosting translation performance of large language models. arXiv preprint arXiv:2309.11674.
- Yang et al. (2023) Wen Yang, Chong Li, Jiajun Zhang, and Chengqing Zong. 2023. Bigtrans: Augmenting large language models with multilingual translation capability over 100 languages. arXiv preprint arXiv:2305.18098.
- Zhang et al. (2023a) Biao Zhang, Barry Haddow, and Alexandra Birch. 2023a. Prompting large language model for machine translation: A case study.
- Zhang et al. (2023b) Shaolei Zhang, Qingkai Fang, Zhuocheng Zhang, Zhengrui Ma, Yan Zhou, Langlin Huang, Mengyu Bu, Shangtong Gui, Yunji Chen, Xilin Chen, et al. 2023b. Bayling: Bridging cross-lingual alignment and instruction following through interactive translation for large language models. arXiv preprint arXiv:2306.10968.
- Zhu et al. (2023) Wenhao Zhu, Hongyi Liu, Qingxiu Dong, Jingjing Xu, Lingpeng Kong, Jiajun Chen, Lei Li, and Shujian Huang. 2023. Multilingual machine translation with large language models: Empirical results and analysis. arXiv preprint arXiv:2304.04675.
Appendix A JSD for Corpora
A.1. JSD of Continued Pre-training Monolingual Data
We use Jensen-Shannon divergence (JSD) as a metric to assess the domain similarity of the HAN dataset. Since there are currently no word segmentation tools available for Hokkien writing systems, the calculation of JSD is a challenging task. The HAN dataset has a certain degree of similarity with Traditional Chinese characters. Therefore, we use the CKIP Traditional Chinese segmentation tool121212https://github.com/ckiplab/ckiptagger to segment the HAN dataset and compute the JSD to analyse the domain similarity of our dataset.
Figure 3 demonstrates that the recitation contest and web-scraped article datasets are more closely aligned. This similarity can be attributed to the fact that both datasets primarily consist of well-structured prose articles, which often explore topics related to local culture, customs and traditions. On the other hand, the content in Hokkien textbooks is composed of verses resembling children’s rhymes, making it more akin to Hokkien song lyrics. Subtitles exhibit a distinctive divergence as they primarily feature colloquial sentence structures, setting them apart from the other two groups.
A.2. JSD of Fine-tuning Parallel Data
In the parallel corpus, we adhere to the CKIP word segmentation system, which is designed for Traditional Chinese. Consequently, we calculate JSD similarity scores based on the Traditional Chinese portion of the text. Figure 4 presents the similarity between the training set, comprising dictionary and technical terms, and the test set, which includes iCorpus-100 and TAT. We observe that the technical terms subset diverges significantly from the others, because it consists solely of terminology, making it less similar to more general sentences. Moreover, there are significant domain differences between the training and test set, indicating that achieving a high performance on the translation with this test set presents a considerable challenge.
Appendix B Details of the GPT-4 evaluation methodology
B.1. Prompt Template
When the target language is in English or Mandarin Chinese:
When the source language is in English or Mandarin Chinese and the target language is in Hokkien:
B.2. Qualitative Analysis of GPT-4 Score Ranges and Examples
In the evaluation of translation scores from GPT-4, we performed a qualitative analysis of a sample of 20 examples from five distinct score intervals, as shown in Table 4. To offer a more comprehensive understanding, we have included examples of translations from Hokkien Han to Mandarin Chinese and English in Tables 8 and 9, respectively.
| Model Translation Results |
100 | |||
70 | |||
50 | |||
30 | |||
0 | |||
Hokkien Han Source: 準備離婚的翁仔姐仔共同攑槌,對準結婚手指殘殘摃落,佇觀眾的噗仔聲中,結婚手指當場歪斜走形。 English Reference: The couple who were preparing for divorce both lifted their hammers and struck hard at their wedding rings. Amidst the applause of the audience, the wedding rings twisted and deformed on the spot. |
| Model Translation Results |
GPT-4 Score |
The couple prepared to divorce and took the hammer together. They hit each other’s wedding rings with force, causing them to deform in front of the audience’s applause. |
90 |
The doll sisters prepared to divorce, and together held the hammer. They hit their wedding ring hard, causing it to bend in front of the audience’s applause. |
60 |
The bride and groom held the hammer together to sign their divorce agreement. They both hit each other’s ring finger hard, causing it to twist in front of the audience’s applause. |
50 |
The doll sisters who are preparing for divorce hold a hammer together, and the ring finger is broken in front of the audience’s applause. On stage, they immediately ran away with their hands crossed. |
20 |
The bride and groom held the hammer together to hit the nails, but they accidentally hit each other’s fingers. In front of the audience, their hands were twisted in a strange way. |
10 |